OpenStudy (itrymath):

4. The following function defines a recursive sequence. f(0) = -5 f(1) = 20 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? (4 points) -5, -20, -65, -200, … -5, 20, -92, 372, … -5, -24, -92, -372, … -5, 20, -65, 200

OpenStudy (itrymath):

@jtm-PA

OpenStudy (itrymath):

we can eliminate A and C

OpenStudy (solomonzelman):

Your \(f(0)=5\) and \(f(1)=20\), so yes, you can eliminate \(A\) and \(C\).

OpenStudy (solomonzelman):

Now, plug in \(f(0)=1\) and \(f(1)=20\) to find \(f(2)\).

OpenStudy (itrymath):

19?

OpenStudy (solomonzelman):

\(f(n)=-4f(n-1)-3f(n-2)\) (for n>1) \(f(2)=-4f(2-1)-3f(2-2)\) \(f(2)=-4f(1)-3f(0)\) \(f(2)=-4(20)-3(5)\)

OpenStudy (solomonzelman):

See what I did?

OpenStudy (itrymath):

-80 - 15 = -95

OpenStudy (solomonzelman):

Yes, that is correct.

OpenStudy (solomonzelman):

Do you have any questions about the procedure though?

OpenStudy (itrymath):

Yes!

OpenStudy (itrymath):

thereis no answer like that

OpenStudy (itrymath):

its not everything is here

OpenStudy (itrymath):

im looking at the original... ill take screenshot

OpenStudy (solomonzelman):

oh damn

OpenStudy (solomonzelman):

\(\bf f(2)=−4(20)−3(\color{red}{-}5)=-65\).

OpenStudy (solomonzelman):

I missed the negative:)

OpenStudy (itrymath):

yyeesss OMGTHATS ITS

OpenStudy (itrymath):

OMG thank youuu

OpenStudy (solomonzelman):

No problem ....

OpenStudy (itrymath):

@SolomonZelman

OpenStudy (itrymath):

whats the slope formula

OpenStudy (solomonzelman):

The formula that gives the slope of a line. (You can look up the equation online ... I lost my famous word doc with copy-paste equations.)